OpenStudy (anonymous):

What is the solution for this system of linear equations? please explain!! y=x-8 y=8x-71

OpenStudy (anonymous):

set them equal to eachother, you will end up with x-8 = 8x -71, solve for x

OpenStudy (anonymous):

then what do you do? im confused?

OpenStudy (anonymous):

start by adding 71 to both sides

OpenStudy (anonymous):

x=63?

OpenStudy (anonymous):

close, divide by the coefficient of the x

OpenStudy (anonymous):

whats the coefficient?

OpenStudy (anonymous):

7

OpenStudy (anonymous):

so I have 8x-71=1x=8 explain to me what I do from here? I subtract 71 from both sides here?

OpenStudy (anonymous):

i would start by subtracting \(x\) from both sides, to put all the terms with \(x\) on the left

OpenStudy (anonymous):

in other words get \(7x-71=-8\)

OpenStudy (anonymous):

solution of a set of equation means  find the point P(x,y) which satisfies all the equations.

OpenStudy (anonymous):

then add 71 to both sides to get \[7x=63\] and finally divide both sides by \&\)

OpenStudy (anonymous):

that, is , finally divide both sides by \(7\)

OpenStudy (anonymous):

so the answer for x=9?

OpenStudy (anonymous):

correct.

OpenStudy (anonymous):

so how would I find y now since I know x?

OpenStudy (anonymous):

you replace \(x\) by \(9\) in either equation

OpenStudy (anonymous):

wait, is y=1?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[y=x-8\] \[y=9-8=1\]

OpenStudy (anonymous):

omg thank you so much! you're amazing!

OpenStudy (anonymous):

set them side by side like this: x-8=8x-71 subtract -8 on both sides and add 8 to both sides now you should get -7x=-63

OpenStudy (anonymous):

yw (if you meant me, if not, then yw from everyone )

OpenStudy (anonymous):

now plug byour answer back into the first equation to get y